OpenStudy (anonymous):

I am confused on 1a. I think we are supposed to use ge(dot)(t) in order to find the body velocity. Is this my ge(dot)(t): (1/T)*[xe(T)-xe(0); ye(T)-ye(0); theta(T)-theta(0)]

OpenStudy (anonymous):

I think our ge from the last homework was: {1.52577; -.547367; -.785398} + (t/T)*{.173704; .542958; 1.309} If I take the derivative of this with respect to t, I get: ge(dot) = (1/T)*{.173704; .542958; 1.309} Am I doing this right?

OpenStudy (anonymous):

Yes that looks right for gedot. You just plugged in the values from last week correct?

OpenStudy (anonymous):

The numbers I used came from partb of problem 1 of last week's hw. The solutions say: "one canidate trajectory connecting the two end effector configurations is: {xe*(t); ye*(t); theta*(t)} = {1.52577; -.547367; -.785398} + (t/T)*{.173704; .542958; 1.309} Is this correct?

OpenStudy (anonymous):

Yes you seem to have taken the derivative properly.

OpenStudy (anonymous):

So, if xi = g^-1(alpha) * (dg(alpha)/dalpha) Then I would simply multiply the inverse of my first ge value by the derivative that I calculated?

OpenStudy (anonymous):

I'm not sure what you mean by your first ge value. You multiply it by the inverse of ge which varies as a function of time. It should look like this: \[\xi_b = \left[\begin{matrix}R^{-1}(\theta_e(t)) & 0 \\ 0 & 1\end{matrix}\right]*gdot_e\]

OpenStudy (anonymous):

What is my ge?

OpenStudy (anonymous):

I can't take the inverse of ge = {1.52577; -.547367; -.785398} + (t/T)*{.173704; .542958; 1.309}

OpenStudy (anonymous):

You don't need to take the inverse of anything. The formula above is the body velocity - it varies as a function of time.

OpenStudy (anonymous):

The theta(t) comes from the trajectory - for any given time t, you can figure out theta and thus the body velocity.

OpenStudy (anonymous):

what is R^-1(theta(t))?

OpenStudy (anonymous):

At this point, I'm not sure if the orginal gedot or ge I posted is correct

OpenStudy (anonymous):

Do you know how to make a rotation matrix \[R(\theta)?\]

OpenStudy (anonymous):

R = [cos(theta), -sin(theta; sin(theta) , cos(theta)]

OpenStudy (anonymous):

Yes so R^-1 is the inverse of that.

OpenStudy (anonymous):

R^-1 = [-cos(theta), -sin(theta); sin(theta), -cos(theta)]

OpenStudy (anonymous):

\[R^{-1}(\theta) = R^T(\theta) = R(-\theta)\]

OpenStudy (anonymous):

So xi = R(-theta) * [(1/T)*[xe(T)-xe(0); ye(T)-ye(0); theta(T)-theta(0)]]

OpenStudy (anonymous):

that would be a 2x2 times a 3x1. That doesn't work. It's what I posted above.

OpenStudy (anonymous):

If gdot = (1/T)*[xe(T)-xe(0); ye(T)-ye(0); theta(T)-theta(0)], how do I get something other than a 3x1 matrix?

OpenStudy (anonymous):

You don't. But you multiply it on the left by \[\left[\begin{matrix}R^{-1}(\theta(t)) & 0 \\ 0 & 1\end{matrix}\right]\]

OpenStudy (anonymous):

Not just R(-theta)

OpenStudy (anonymous):

So, xi = [cos(-theta(t)), -sin(-theta(t)), 0; sin(-theta(t)), cos(-theta(t)), 0; 0 , 0, 1] * [(1/T)*[xe(T)-xe(0); ye(T)-ye(0); theta(T)-theta(0)]]

OpenStudy (anonymous):

Yes, that xi_body. I gotta go. Good luck

OpenStudy (anonymous):

Thanks for the help.

jaynator495 (jaynator495):

Welcome To OpenStudy! Here you will find great helpers and friends, a community of students that help students! We hope you enjoy the experience!